OpenStudy (anonymous):

what are solutions to this equation 2x^2+6x=20 the options r -5,-10,-2,1,2,20

OpenStudy (mathmale):

I see immediately that we could simplify the work involved by dividing every term of this equation by 2. Would you do that , please?

OpenStudy (solomonzelman):

2x^2+6x=20 2x^2+6x-20 =0 2(x^2+3x-10) =0 x^2+3x-10 =0 (think of 5 and -2 )

OpenStudy (mathmale):

As a next step, please consolidate all of the terms on the left side. To do this, please subtract 10 from both sides. Type out the result here.

OpenStudy (acxbox22):

divide everything by 2 complete the square

OpenStudy (anonymous):

Hint: its two of the options that are given to you

OpenStudy (acxbox22):

nice hint

OpenStudy (anonymous):

Thank you

OpenStudy (acxbox22):

x^2+3x=10 x^2+3x+2.25=12.25 (x+1.5)^2=12.25 find squareroot of each \[x+1.5=\pm3.5\]

OpenStudy (acxbox22):

x=3.5-1.5 x=-3.5-1.5 x=2,-5